OpenStudy (anonymous):

Multiply (x + 5)(x - 6). A) x2 - x - 30 B) x2 + x - 30 C) x2 - 11x - 30 D) x2 + 11x - 30

OpenStudy (anonymous):

x*x=x^2 x*-6=-6x x*5=5x 5*-6=-30 Combine like terms. x^2-x-30

OpenStudy (anonymous):

Use FOIL. F-first O-outer I-inner L-last